# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v3

@@ 4.5. RQ4: Parameter Sensitivity
To investigate the impact of lookahead length $L$ on decoding performance, we conduct experiments on HumanEval+ using DS-1.3B as the backbone model. We vary $L$ from 2 to 9 and report pass@1 results. Figure 6 illustrates the trend.
@@
The results show that pass@1 increases steadily with larger $L$ up to $L=5$, at which point it reaches the peak performance. Beyond $L=5$, the performance declines slightly and then plateaus. This behavior can be explained by the trade-off between local guidance and global noise: shorter lookaheads may fail to capture sufficient future context, while excessively long lookaheads introduce noise and dilute the relevance of local signals. Importantly, across all values of $L$, AdaDec consistently outperforms the baseline decoding strategies, demonstrating the robustness of our approach to the choice of lookahead length.
Summary: The lookahead length $L$ exhibits a non-monotonic effect on performance: moderate values (e.g., $L=5$) maximize the benefits of lookahead guidance, while excessively small or large values reduce effectiveness. Notably, AdaDec consistently surpasses baseline performance regardless of the choice of $L$.
@@ 5.1. Limitations
Potential Bias from Training Data. Although we carefully avoided direct overlap between training and evaluation datasets, our threshold learning still relies on synthetic data generated from a specific corpus. If the distribution of training data diverges significantly from the deployment scenario, the learned thresholds may fail to transfer robustly. Future work could explore continual adaptation or online threshold adjustment.
@@ 6. Threats to Validity
External Validity. We focus on open-source models (DeepSeek, StableCode, Qwen3) and three benchmarks (HumanEval+, MBPP+, DevEval). Results may not fully generalize to proprietary models (e.g., GPT-4, Codex) or to broader real-world tasks. Future work should extend evaluation to more diverse datasets and model families.
@@ 7.1. LLM-based Code Generation
To evaluate their effectiveness, benchmarks like HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), and DS-1000 (Lai et al., 2022) focus on function-level generation, while ClassEval (Du et al., 2023) extends evaluation to class-level tasks. More recently, CoderEval (Zhang et al., 2024) and DevEval (Li et al., 2024) target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.

### v4

@@ 4.5. RQ4: Parameter Sensitivity
To investigate the impact of key hyperparameters on decoding performance, we conduct sensitivity analyses on the lookahead length $L$ and the learned entropy threshold $\tau^{LM}$. All experiments in this section use DS-1.3B as the backbone model evaluated on HumanEval+.
Sensitivity to Lookahead Length ($L$). We vary $L$ from 2 to 9 and report Pass@1 results. Figure 6 illustrates the trend.
@@
The results show that Pass@1 increases steadily with larger $L$ up to $L=5$, at which point it reaches the peak performance of 65.24%. Beyond $L=5$, the performance does not continue to rise; instead, it exhibits slight fluctuations, hovering between 63.41% and 64.63%. This behavior reflects the trade-off between local guidance and global noise: while shorter lookaheads may fail to capture sufficient future context, excessively long lookaheads ($L>5$) may introduce noise from uncertainty accumulation, diluting the relevance of immediate local signals.
Crucially, across all values of $L$, AdaDec consistently outperforms the Greedy baseline (62.20%). Even at the minimal setting of $L=2$, AdaDec achieves 62.80%, reducing to a safe margin of +0.60%. This demonstrates that our approach is robust to hyperparameter choices and provides stable improvements.
Sensitivity to Threshold Variations ($\tau^{LM}$). As illustrated in Table 7, we evaluate the Pass@1 accuracy by shifting the optimal learned threshold ($\tau^{LM}\approx 0.9850$) across various intervals. AdaDec demonstrates high robustness to minor fluctuations. Within the $\tau^{LM}-0.1,\tau^{LM}+0.1]$ range, Pass@1 remains highly stable and consistently outperforms the greedy baseline.
Table 7. Sensitivity of AdaDec performance (Pass@1) to entropy threshold ($\tau$) variations on HumanEval+. Evaluated with DS-1.3B. Learned denotes the optimal threshold derived from our approach.
Setting Threshold $\tau$ Pass@1 (%) Pause Rate (%)
Baseline (Greedy) 62.20
$\tau^{LM}-0.50$ 0.4850 63.41 7.82
$\tau^{LM}-0.10$ 0.8850 65.24 3.46
$\tau^{LM}-0.05$ 0.9350 64.63 3.10
$\tau^{LM}$ (Learned) 0.9850 65.24 2.81
$\tau^{LM}+0.05$ 1.0350 65.24 2.52
$\tau^{LM}+0.10$ 1.0850 64.63 2.16
$\tau^{LM}+0.50$ 1.4850 62.80 0.83
Visible performance degradation occurs only under drastic shifts ($\pm 0.5$): excessively low thresholds introduce noise via over-intervention, while excessively high thresholds become too conservative. This confirms that $\tau^{LM}$ is not a fragile hyperparameter requiring extreme precision.
Summary: Both key hyperparameters exhibit high robustness. The lookahead length $L$ shows a non-monotonic effect, with moderate values (e.g., $L=5$) maximizing the benefits of lookahead guidance. The learned threshold $\tau^{LM}$ is highly stable against minor fluctuations within $\pm 0.1$. Notably, AdaDec consistently surpasses baseline performance across varying hyperparameter settings.
@@ 5. Discussion
5.1. Generalizability and Transferability of the Learned Threshold ($\tau^{LM}$)
We conduct a quantitative analysis using DS-1.3B to evaluate the robustness of the learned entropy threshold ($\tau^{LM}$) across domain and prompt shifts, and provide guidance on its practical application.
Robustness to Domain and Prompt Shifts. We evaluate the transferability of $\tau^{LM}$ to library-heavy domains and varying prompt templates. We compute the optimal thresholds for two specific data science scenarios: a data science subset filtered from BigCodeBench (666 problems) and the DS-1000 benchmark (Lai et al., 2022) (1000 problems, featuring drastically different prompt styles and task requirements). The optimal thresholds for these two datasets are $0.9740$ ($\Delta=0.0110$) and $0.9178$ ($\Delta=0.0672$), respectively. Since our parameter sensitivity analysis (Section 4.5) demonstrates that deviations within $\pm 0.1$ do not significantly impact performance, a globally learned $\tau^{LM}$ remains highly effective without the need for task-specific tuning.
Guidance on Relearning $\tau^{LM}$. These findings indicate that $\tau^{LM}$ reflects the intrinsic uncertainty calibration of the base LLM rather than task-specific artifacts. Therefore, relearning $\tau^{LM}$ is generally unnecessary for different standard coding domains or prompt styles. Retraining is only recommended when the base model’s confidence distribution changes fundamentally, such as after Supervised Fine-Tuning (SFT) or alignment techniques.
@@ 5.2. Limitations
Fixed Lookahead Depth. Currently, AdaDec utilizes a fixed lookahead length (e.g., $L=5$) across all benchmarks. While this serves as a reasonable global setting, repository-level tasks (e.g., DevEval) often require longer-range reasoning, and a short lookahead window may under-express these dependencies. Future research could explore adaptive, context-aware lookahead depths that dynamically adjust to accommodate both short-term syntactic constraints and long-term semantic reasoning demands.
@@ 6. Threats to Validity
External Validity. We focus on open-source models (DeepSeek, StableCode, Qwen) and three benchmarks (HumanEval+, MBPP+, DevEval). Results may not fully generalize to proprietary models (e.g., GPT-4, Codex) or to broader real-world tasks. Future work should extend evaluation to more diverse datasets and model families.
@@ 7.1. LLM-based Code Generation
To evaluate their effectiveness, benchmarks like HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and DS-1000 (Lai et al., 2022) focus on function-level generation, while ClassEval (Du et al., 2023) extends evaluation to class-level tasks. More recently, CoderEval (Zhang et al., 2024) and DevEval (Li et al., 2024) target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.